0.500 g of a sample which contains copper is analysed iodometrically. Copper (ll) is reduced to copper (l) by the iodide. What is the percent copper in the sample if 20.00 cm^3 of 0.1M sodium thiosulphate solution is required for the titration of the liberated iodine?


1
Expert's answer
2021-03-22T08:03:19-0400

Volume of solution = 20cm3 = 0.02L

molarity = 0.1M

no. Of moles = molarity × volume

moles = 0.02 ×0.1 = 0.002

Mass of Cooper in solution = 0.002×63.546

= 0.127 g

% copper = (0.127 / 0.500) × 100 = 25.4%